OpenStudy (anonymous):

given that angle A and angle B are complementary if angle A (2x-3) and angle B =(3x=8) find angle A

OpenStudy (anonymous):

SUM OF A AND B IS 90 2X-3+3X+8=90 5X=85 X=17 A=30

OpenStudy (anonymous):

2x-3+3x+8=90 5x+5=90 5x=90-5 5x=85 x=17 angle A is 2x-3 =2(17)-3 =34-3 =31 degrees

OpenStudy (anonymous):

YEP SORRY IT IS 31

OpenStudy (anonymous):

it is okay..